OpenStudy (anonymous):

implciitely differentiate x^2+y^2=L^2 is it 2xdx/dx +2ydy/dt=0 becase L is a function of nothing??

OpenStudy (anonymous):

What are you differentiating with respect to? x? And what does L represent here? Is it a constant?

OpenStudy (anonymous):

well L is a fixed given L=25ft where as x fixes is given at x=7, y=sqroot 25^2-7^2=24 x and y are both changing with respect to t (its a related rates problem) 25ft represent the size "a 25 ft ladder"

OpenStudy (anonymous):

Okay that sounds right then. So you get 2x+2yy'=0, y'=-x/y

OpenStudy (anonymous):

The Total Derivative of x^2+y^2=L^2 is:\[2 x dx+2 y dy=2 L dL \]if L is a constant, then dL is zero and the RHS is zero.\[2 x dx+2 y dy=0 \]

OpenStudy (anonymous):

ty